OpenStudy (anonymous):

(sqrt2z+6)+(sqrtz+4)-(sqrt8z+9)=0

OpenStudy (ajprincess):

\((\sqrt{2z}+6)+(\sqrt{z}+4)-(\sqrt{8z}+9)=0\) Is this the question?

OpenStudy (anonymous):

yes thats the question to be solved. thnx

OpenStudy (ajprincess):

\(\sqrt{8z}=\sqrt{2*2*2*z}\) \(=2\sqrt{2z}\) \((\sqrt{2z}+6)+(\sqrt{z}+4)-(2\sqrt{2z}+9)=0\) \(\sqrt{2z}+6+\sqrt{z}+4-2\sqrt{2z}-9=0\) This simplifies to \(1+\sqrt{z}-\sqrt{2z}=0\) \(\sqrt{2z}-\sqrt{z}=1\) Factor out \(\sqrt{z}\). Then u will get \(\sqrt{z}(\sqrt{2}-1)=1\) Square both sides and solve for z. Can u do it @yayimoakamie?

OpenStudy (anonymous):

ok thnx let me give it a try

OpenStudy (ajprincess):

sure:)

OpenStudy (anonymous):

is the final answer z=1

OpenStudy (ajprincess):

Hmm no.

OpenStudy (ajprincess):

Did u round off ur answr?

OpenStudy (ajprincess):

@yayimoakamie

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i got it wrong then

OpenStudy (ajprincess):

yup. \((\sqrt2-1)^2=(\sqrt2)^2-2*\sqrt2*1+1^2\) \(=2-2\sqrt2+1\) \(=3-2\sqrt2\) \((\sqrt z)^2=z\) \(z(3-2\sqrt2)=1\) \(z=\LARGE\frac{1}{3-2\sqrt2}\) Getting this @yayimoakamie?

OpenStudy (anonymous):

Thnx Ajprincess for your help appreciated

OpenStudy (ajprincess):

yw:) do u understand it?

OpenStudy (anonymous):

honestly i am lost

OpenStudy (ajprincess):

Tell me where u r lost?

OpenStudy (anonymous):

why is the square root not affecting the constants but just the variables

OpenStudy (ajprincess):

Sorry. I dnt get ur question.

OpenStudy (ajprincess):

can u plz show where it is nt being affected?

OpenStudy (ajprincess):

@yayimoakamie

OpenStudy (anonymous):

sorry what i mean is d constants 6+4+9 were not affected by the root

OpenStudy (ajprincess):

in ur original question square root is for only z right? That is y they are not being affected.

OpenStudy (ajprincess):

when u solve an equation u solve for the variable. the variable should be alone without square root or anything.